# On manifolds with almost non-negative Ricci curvature and integrally-positive $k^{th}$-scalar curvature

### v2

@@ 3.2 First geometric applications in almost non-negative Ricci and integrally-positive scalar curvature
Let $(X,\mathsf{d},x_{\infty})$ be a pGH limit of a (non-relabeled) subsequence of $(M_{i},g_{i},p_{i})$. This limit exists thanks to Remark 3.9. Let $x\in B_{1}(x_{\infty})\setminus x_{\infty}$ be a point where $\mathbb{R}^{d}$ is a tangent space of $X$. We claim that $d\leq k-1$. If this holds, then a rescaling argument implies the same dimensional bound on tangent spaces for every $x\neq x_{\infty}$.
Let $l:=\mathsf{d}(x_{\infty},x)$, and let $R\gg 1$. Fix $\epsilon>0$ and let $0<s\ll(l/2R)$ be small enough so that
  • $B^{i}_{s}(p^{\prime}_{i})\subset B^{i}_{1}(p_{i})$ for $i$ large enough.
  • $(X,\mathsf{d}/s,x)$ is $\epsilon$-close in pGH sense to $\mathbb{R}^{d}$.
Then, there exists points $p^{\prime}_{i}\in M_{i}\setminus B^{i}_{l/2}(p_{i})$ converging to $x$ such that, for $i$ large enough, $(M_{i},g_{i}/s^{2},p^{\prime}_{i})$ is $2\epsilon$-close in pGH sense to $\mathbb{R}^{d}$.
Moreover, if $y\in B^{i}_{Rs}(p^{\prime}_{i})$, it holds $\mathsf{d}_{i}(y,p_{i})\geq l/2-Rs>0$, which is equivalent to
$\mathsf{d}(y,p)\geq r_{i}^{2}(l/2-Rs)>0.$ (23)
Hence,
$\mathsf{Ric}_{M_{i}}(y)\wedge 0\geq r_{i}^{2}\mathsf{Ric}_{M}(y)\wedge 0\geq(l/2-Rs)^{-2}\mathsf{d}(p,y)^{2}\mathsf{Ric}_{M}(y)\wedge 0.$
Since $M$ has asymptotically non-negative Ricci curvature, combining (23) with the fact that $r_{i}\uparrow+\infty$, we deduce that, for every $\delta>0$, it holds
$\mathsf{Ric}_{M_{i}}\geq-\delta\quad\text{on }B^{i}_{Rs}(p^{\prime}_{i})\text{for }i\text{ large enough}.$ (24)
We claim that, for $i$ large enough, we also have
@@
If this is true, by choosing $\epsilon>0$ small enough and combining with (24), we deduce that $d\leq k-1$ by Theorem 3.12.
@@
Using Lemma 3.10 in the last inequality, we find a constant $c(s,n)>0$ such that
@@
Let $L>0$ and $\alpha\in(0,2)$ be fixed. Another immediate consequence of Theorem 3.14 is that if a manifold $(M,g)$ with asymptotically non-negative Ricci curvature has $\mathsf{Ric}_{M}\geq f$ for a function $f$ such that
@@
then $M$ is compact. A similar result was obtained in [CGT, Theorem 4.8], where also the case $\alpha=2$ was addressed, under a pointwise (rather than integral) bound on $\mathsf{Ric}_{M}$.
@@
Proof of 1: Let $(\tilde{M},\tilde{g},\tilde{p})$ be a covering of $M$, and let $\pi:\tilde{M}\to M$ be the covering map. We claim that $\tilde{M}$ has a tangent cone at infinity of dimension $\leq k-1$.
@@
As a consequence, by item 1 of Theorem 3.14 and Remark 3.15, $\tilde{M}$ has a tangent cone at infinity of dimension $\leq k-1$. The statement now follows by repeating the argument of [zhu2024twodimensionvanishingsplittingpositive, Theorem 1.3], where instead of using [zhu2024twodimensionvanishingsplittingpositive, Theorem 1.6] one uses that $\tilde{M}$ has a tangent cone at infinity of dimension $\leq k-1$.
@@
Remark 3.19.
In the proof of the previous result, one cannot weaken condition (27) to
$\limsup_{x\to+\infty}\mathsf{d}(p,x)^{2-\alpha}\mathsf{R}_{k}(x)>0.$ (29)
Indeed, (29) does not pass to covering spaces. To see this, let $k=3$ and consider a product ×$\mathbb{R}\times S^{1}\times S^{2}$ with the standard metric. We can now perturb the metric in a compact set so that it has negative $\mathsf{R}_{3}$ at a point. The universal cover of this perturbed space will then have a sequence of points with negative $\mathsf{R}_{3}$ escaping to infinity. A similar argument holds for condition (28).
@@
and $\mathrm{b}_{1}(M_{j})\geq k$. Up to a subsequence, $M_{j}$ GH-converge to a metric space $(X,\mathsf{d})$. Thanks to the non-collapsing assumption, the metric measure space $(X,\mathsf{d},\mathsf{H}^{n})$ is an $\mathrm{RCD}(0,n)$ space while, thanks to Proposition 2.16, it holds $\mathrm{b}_{1}(X)\geq k$. By Proposition 2.20, the universal cover $(\tilde{X},\tilde{\mathsf{d}})$ of $X$ splits $\mathbb{R}^{k}$ isometrically.
Let $\tilde{p}\in\tilde{X}$. By Proposition 2.21, the pointed metric space $(\tilde{X},\tilde{\mathsf{d}},\tilde{p})$ is the pGH-limit of a sequence $(\tilde{M}_{j},\tilde{g}_{j},\tilde{p}_{j})$, where each $\tilde{M}_{j}$ covers $M_{j}$. These covering spaces satisfy
@@ 4 Thin metric spaces
The goal of this section is to show that thin metric spaces (see Definition 4.2, after [TrianglesKapo]) are contained in a neighbourhood of controlled width of an isometrically embedded $1$-dimensional manifold (see Theorem 4.16). This will be a key step to prove Theorem 1.2.
In [TrianglesKapo, Section 5], it is shown that a non-compact thin metric space is contained in a finite neighourhood of a ray or a line. Comparing with [TrianglesKapo, Section 5], Theorem 4.16 treats the compact case as well, and provides explicit uniform bounds on the width of the aforementioned neighbourhood. These bounds are needed for the applications in the subsequent sections and, in particular, to prove Theorem 1.2. On the other hand, unlike [TrianglesKapo, Section 5], we only consider metric spaces $(X,\mathsf{d})$ that are proper, complete, separable, and geodesic.
@@
Let $R,D>0$ with $R\geq 20D$. We say that the metric space $(X,\mathsf{d})$ is $(R,D)$-thin if for every segment $r$ of length greater than $2R$ in $X$, and for every $t\in(R,\mathrm{L}(r)-R)$ and $x\in\pi_{r}^{-1}(r(t))$, it holds $\mathsf{d}(x,r)<D$.
In [TrianglesKapo], a metric space is defined to be $R$-thin if it is $(R,R)$-thin according to Definition 4.2. We consider the extra parameter $D>0$ and we restrict ourselves to the case $R\geq 20D$ as this will be sufficient for the subsequent applications.
@@
Let $(X,\mathsf{d})$ be $(R,D)$-thin (see Definition 4.2) and let $r$ be a ray in $X$. If a point $x\in X$ has a projection on $r$ which lies in $(r(R),r(\infty))$, then $\mathsf{d}(x,r)\leq D$. This fact follows by applying Definition 4.2 to the segments $[r(0),r(T)]$ for $T>2R$.
@@
Let $(X,\mathsf{d})$ be a metric space and let $C_{1}\subset C_{2}\subset X$ be closed sets. The set
$K:=\{x\in X:x\text{ has a projection on }C_{2}\text{ lying in }C_{1}\}$
@@
$I:=\{s\in[0,b]:\alpha(s)\text{ has a projection on }C_{2}\text{ lying in }C_{1}\}$
@@
    Let $l$ be a segment with $\mathrm{L}(l)>2R$. Let $t\in(R,\mathrm{L}(l)-R)$, and let $u\in[0,R]\cup[\mathrm{L}(l)-R,\mathrm{L}(l)]$. If $l(t)\in\pi_{l}(\alpha(0))$ and $l(u)\in\pi_{l}(\alpha(\mathrm{L}(\alpha)))$, then the set $\pi_{l}(\alpha)$ contains a $3D$-net of either $[l(R),l(t)]$ or $[l(t),l(\mathrm{L}(l)-R)]$.
@@
    Let $l$ be a segment with $\mathrm{L}(l)>2R$. Let $t,u\in(R,\mathrm{L}(l)-R)$ with $t<u$. If $l(t)\in\pi_{l}(\alpha(0))$ and $l(u)\in\pi_{l}(\alpha(\mathrm{L}(\alpha)))$, then $\pi_{l}(\alpha)$ contains a $3D$-net of either $[l(t),l(u)]$ or $[l(R),l(t)]\cup[l(u),l(\mathrm{L}(l)-R)]$.
@@
$\tilde{s}:=\max\Big{\{}s\in[0,$ $\mathrm{L}(\alpha)]:\alpha(s)\text{ has a projection on }l$
[…]
In order to show that $\tilde{s}$ is well defined, it is sufficient to observe that the set appearing in its definition is:
@@
We claim that $\tilde{s}<\mathrm{L}(\alpha)$. If this is not the case, then, combining with (31), $\alpha(\mathrm{L}(\alpha))$ admits a projection on $l$ contained in $[l(x_{1}+3D),l(x_{2}-3D)]$. Since $X$ is $(R,D)$-thin, $\alpha(\mathrm{L}(\alpha))$ is $D$-close to $[l(x_{1}+3D),l(x_{2}-3D)]$, so that it cannot have a projection on $l$ given by $l(u)$ for $u\in[0,R]\cup[\mathrm{L}(l)-R,\mathrm{L}(l)]$, a contradiction. This shows that $\tilde{s}<\mathrm{L}(\alpha)$.
Hence, for every $\epsilon>0$, there is a projection $p_{\epsilon}$ of $\alpha(\tilde{s}+\epsilon)$ on $l$. Let $p$ be a projection of $\alpha(\tilde{s})$ on $l$ lying in $[l(x_{1}),l(x_{2})]$. Since $X$ is $(R,D)$-thin, it holds $\mathsf{d}(p,\alpha(\tilde{s}))<D$. It follows that $\mathsf{d}(p_{\epsilon},\alpha(\tilde{s}+\epsilon))<D+\epsilon$. Combining these facts, it holds
@@
We now consider item 3. Since the image of $\alpha$ is compact, $\pi_{l}(\alpha)$ is a bounded subset of $l$. Hence, $\pi_{l}(\alpha)\subset[r(0),r(T)]$ for some $T>0$. We consider now the segment $l:=[r(0),r(T+2R)]$. It is easy to check that $\pi_{l}(\alpha)=\pi_{r}(\alpha)$. By item 1, $\pi_{l}(\alpha)$ contains a $3D$-net of either $[l(R),l(t)]$ or $[l(t),l(\mathrm{L}(l)-R)]$. This last case cannot happen as $[l(t),l(\mathrm{L}(l)-R)]=[r(t),r(T+R)]$ and $\pi_{l}(\alpha)\subset[r(0),r(T)]$. Hence, $\pi_{l}(\alpha)$ (which again coincides with $\pi_{r}(\alpha)$) contains a $3D$-net of $[l(R),l(t)]=[r(R),r(t)]$. ∎
@@
The next proposition shows that an $(R,D)$-thin non-compact metric space is contained in a neighbourhood of controlled width of a ray or a line. We recall that a similar result is proved in [TrianglesKapo, Section 5]. The main difference is that Proposition 4.10 below also gives an explicit bound on the width of the aforementioned neighbourhood.
@@
We now prove the statement of the lemma. If $X$ is contained in the $10R$-neighbourhood of $r$, there is nothing to prove. Otherwise, we claim that there exists an element $y\in X$ at maximal distance from $r$. Let $y_{i}\in X$ be a sequence maximizing the distance from $r$. We can assume that $9R\leq\mathsf{d}(y_{i},r)\leq D^{\prime}(X)$. Since $X$ is $(R,D)$-thin, all projections of $y_{i}$ on $r$ are contained in $[r(0),r(R)]$. It follows that the sequence $\{y_{i}\}_{i\in\mathbb{N}}$ is precompact, so that there exists a limit point $y\in X$ at maximal distance from $r$.
Let $r_{y}$ be the ray which is the limit of the segments joining $y$ and $r(t_{i})$ for a sequence $t_{i}\uparrow+\infty$. By Lemma 4.9, for every $t\geq 0$, it holds $\mathsf{d}(r(t),r_{y})\leq 2R$. Since $\mathsf{d}(y,r)\geq 10R$, it follows that all projections of $r$ on $r_{y}$ are contained in $[r_{y}(5R),r_{y}(+\infty))$.
Assume now by contradiction that there exists $y^{\prime}\in X$ such that $\mathsf{d}(y^{\prime},r_{y})\geq 10R$. Let $\gamma$ be a segment realizing the distance between $y^{\prime}$ and $r$. Let $r(t_{\gamma})$ be the endpoint of $\gamma$. Let $r_{y}(t_{\gamma}^{y})$ be a projection on $r_{y}$ of $r(t_{\gamma})$. By our previous remarks, $t_{\gamma}^{y}\geq 5R$. By Lemma 4.5, $\pi_{r_{y}}(\gamma)$ contains a $3D$-neighbourhood of $[r_{y}(R),r_{y}(5R)]$. Hence, $\gamma$ contains a point $p_{1}$ which is $7D$-close to $r_{y}(R)$. Recall that we also have $\mathsf{d}(y^{\prime},r_{y})\geq 10R$. Hence,
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space. Let $l$ be a segment in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $\bar{l}$ be a segment starting from $l(0)$ with $\pi_{l}(\bar{l}(\mathrm{L}(\bar{l})))\subset[l(0),l(R)]\cup[l(\mathrm{L}(l)-R),l(\mathrm{L}(l)]$. If $\bar{l}$ has a projection on $l$ lying in $[l(3R),l(\mathrm{L}(l)-3R)]$, then $\pi_{l}(\bar{l})$ contains a $3D$-net of $[l(R),l(\mathrm{L}(l)-R)]$.
@@
Let $t_{0}\in[0,\mathrm{L}(\bar{l})]$ be such that $\bar{l}(t_{0})$ has a projection on $l$ given by $l(t)$ for $t\in[3R,\mathrm{L}(l)-3R]$. Let $\bar{l}_{1}:=\bar{l}_{|[0,t_{0}]}$ and $\bar{l}_{2}:=\bar{l}_{|[t_{0},\mathrm{L}(\bar{l})]}$.
Hence, both $\bar{l}_{1}$ and $\bar{l}_{2}$ have an endpoint whose projection on $l$ lies in $[l(0),l(R)]\cup[l(\mathrm{L}(l)-R),l(\mathrm{L}(l)]$.
By item 1 of Lemma 4.5, $\pi_{l}(\bar{l}_{1})$ contains a $3D$-net of either $[l(R),l(t)]$ or $[l(t),l(\mathrm{L}(l)-R)]$, and the same holds for $\pi_{l}(\bar{l}_{2})$.
@@
Since $\bar{l}(t_{0})$ is $D$-close to a point $l(t)$ in $[l(3R),l(\mathrm{L}(l)-3R)]$, it follows
@@
By the same argument, it is not possible that both $\pi_{l}(\bar{l}_{1})$ and $\pi_{l}(\bar{l}_{2})$ contain a $3D$-net of $[l(t),l(\mathrm{L}(l)-R)]$, concluding the proof. ∎
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space. Let $l$ be a segment in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $p\in X$ be a point with $\mathsf{d}(p,l)\geq 10R$ and let $\bar{l}$ be a segment connecting $p$ to $l(0)$. If $\bar{l}$ has a projection on $l$ lying in $[l(3R),l(\mathrm{L}(l)-3R)]$, then $\mathrm{L}(\bar{l})>\mathrm{L}(l)$.
@@
We now consider the case $b_{1}\in\{\gamma_{1},\gamma_{2}\}$ and $b_{2}=l$. If the statement fails, $l$ has a projection on $b_{1}$ lying in $(b_{1}(3R),b_{1}(\mathrm{L}(b_{1})-3R))$. Consider the orientation of $l$ such that $l(0)\in b_{1}$. Then, $\pi_{b_{1}}(l(\mathrm{L}(l)))\subset[b_{1}(0),b_{1}(R)]\cup[b_{1}(\mathrm{L}(b_{1})-R),b_{1}(\mathrm{L}(b_{1})]$ since otherwise, using that $X$ is $(R,D)$-thin, it follows that $l$ is shorter than $b_{1}$. Hence, by Lemma 4.11, there exists $s>0$ such that $l(s)$ is $R+7D$-close to $p$. This contradicts that $\mathsf{d}(p,l)\geq 10R$.
@@
Assume by contradiction that $\gamma_{1}(t)$, with $t\in(3R,\mathrm{L}(\gamma_{1})-3R)$ is a projection of $\gamma_{2}$ on $\gamma_{1}$. By Lemma 4.11, there exists $s>0$ such that $\gamma_{2}(s)$ is $R+7D$-close to $\gamma_{1}(0)=l(\mathrm{L}(l))$. It follows that
@@
By the previous lemma, a projection of $b_{1}(0)$ on $b_{2}$ lies in $[b_{2}(3R),b_{2}(\mathrm{L}(b_{2})-3R]$. If such projection lies in $[b_{2}(R),b_{2}(3R)]$ or $[b_{2}(\mathrm{L}(b_{2})-3R),b_{2}(\mathrm{L}(b_{2})-R)]$, then $\gamma$ has a point which is $5R$-close to either $b_{2}(3R)$ or $b_{2}(\mathrm{L}(b_{2})-3R)$ (because $X$ is $(R,D)$-thin), otherwise the same conclusion follows by Lemma 4.5. ∎

### v3

@@ 3.2 First geometric applications in almost non-negative Ricci and integrally-positive scalar curvature
Let $(X,\mathsf{d},x_{\infty})$ be a pGH limit of a (non-relabeled) subsequence of $(M_{i},g_{i},p_{i})$. Let $x\in B_{1}(x_{\infty})$ be a point where $\mathbb{R}^{d}$ is a tangent space of $X$. We claim that $d\leq k-1$. If this holds, then a rescaling argument implies the same dimensional bound on tangent spaces for every $x\in X$.
Fix $\epsilon>0$, and let $s>0$ be small enough so that $(X,\mathsf{d}/s,x)$ is $\epsilon$-close in pGH sense to $\mathbb{R}^{d}$. Then, there exists points $p^{\prime}_{i}\in M_{i}$ converging to $x$ such that, for $i$ large enough, $(M_{i},g_{i}/s^{2},p^{\prime}_{i})$ is $2\epsilon$-close in pGH sense to $\mathbb{R}^{d}$. Modulo reducing $s$, we may assume that $B^{i}_{s}(p^{\prime}_{i})\subset B^{i}_{1}(p_{i})$ for $i$ large enough.
We claim that, for $i$ large enough, it holds
@@
If this is true, by choosing $\epsilon>0$ small enough, we deduce that $d\leq k-1$ by Theorem 1.1.
@@
By Bishop-Gromov’s inequality, we conclude that for a positive constant $c(s,n)>0$ it holds
@@
Let $L>0$ and $\alpha\in(0,2)$ be fixed. Another immediate consequence of Theorem 3.9 is that if a manifold $(M,g)$ with non-negative Ricci curvature has $\mathsf{Ric}_{M}\geq f$ for a function $f$ such that
@@
then $M$ is compact. A similar result was obtained in [13, Theorem 4.8], where also the case $\alpha=2$ was addressed, under a pointwise (rather than integral) bound on $\mathsf{Ric}_{M}$.
@@
[…]
@@
As a consequence, by item 1 of Theorem 3.9 and Remark 3.10, $\tilde{M}$ has a tangent cone at infinity of dimension $\leq k-1$. The statement now follows by repeating the argument of [64, Theorem 1.3], where instead of using [64, Theorem 1.6] one uses that $\tilde{M}$ has a tangent cone at infinity of dimension $\leq k-1$.
@@
[…]
@@ 4 Thin metric spaces
The goal of this section is to show that thin metric spaces (see Definition 4.2, after [34]) are contained in a neighbourhood of controlled width of an isometrically embedded $1$-dimensional manifold (see Theorem 4.16). This will be a key step to prove Theorem 1.2.
In [34, Section 5], it is shown that a non-compact thin metric space is contained in a finite neighbourhood of a ray or a line. Comparing with [34, Section 5], Theorem 4.16 treats the compact case as well, and provides explicit uniform bounds on the width of the aforementioned neighbourhood. These bounds are needed for the applications in the subsequent sections and, in particular, to prove Theorem 1.2. On the other hand, unlike [34, Section 5], we only consider metric spaces $(X,\mathsf{d})$ that are proper, separable, and geodesic.
@@
Let $R,D>0$ with $R\geq 20D$. Let $(X,\mathsf{d})$ be a proper, geodesic and separable metric space. We say that $(X,\mathsf{d})$ is $(R,D)$-thin if for every segment $r$ of length greater than $2R$ in $X$, and for every $t\in(R,\mathrm{L}(r)-R)$ and $x\in\pi_{r}^{-1}(r(t))$, it holds $\mathsf{d}(x,r)<D$.
In [34], a metric space is defined to be $R$-thin if it is $(R,R)$-thin according to Definition 4.2. We consider the extra parameter $D>0$ and we restrict ourselves to the case $R\geq 20D$ as this will be sufficient for the subsequent applications.
@@
Let $(X,\mathsf{d})$ be $(R,D)$-thin (see Definition 4.2) and let $r$ be a ray in $X$. If a point $x\in X$ satisfies $\pi_{r}(x)\cap(r(R),r(\infty))\neq\emptyset$, then $\mathsf{d}(x,r)\leq D$. This fact follows by applying Definition 4.2 to the segments $[r(0),r(T)]$ for $T>2R$.
@@
Let $(X,\mathsf{d})$ be a geodesic metric space and let $C_{1}\subset C_{2}\subset X$ be closed sets. The set
$K:=\{x\in X:\pi_{C_{2}}(x)\cap C_{1}\neq\emptyset\}$
@@
$I:=\{s\in[0,b]:\pi_{C_{2}}(\alpha(s))\cap C_{1}\neq\emptyset\}$
@@
    Let $l$ be a segment with $\mathrm{L}(l)>2R$. Let $t\in(R,\mathrm{L}(l)-R)$, and let $u\in[0,R]\cup[\mathrm{L}(l)-R,\mathrm{L}(l)]$. If $l(t)\in\pi_{l}(\alpha(0))$ and $l(u)\in\pi_{l}(\alpha(\mathrm{L}(\alpha)))$, then the set $\pi_{l}(\alpha)$ contains a $3D$-net of either $[l(R),l(t)]$ or $l(t),l(\mathrm{L}(l)-R)]$.
@@
    Let $l$ be a segment with $\mathrm{L}(l)>2R$. Let $t,u\in(R,\mathrm{L}(l)-R)$ with $t<u$. If $l(t)\in\pi_{l}(\alpha(0))$ and $l(u)\in\pi_{l}(\alpha(\mathrm{L}(\alpha)))$, then $\pi_{l}(\alpha)$ contains a $3D$-net of either $[l(t),l(u)]$ or $l(R),l(t)]\cup[l(u),l(\mathrm{L}(l)-R)]$.
@@
$\tilde{s}:=\max\Big\{s\in[0,$ $\mathrm{L}(\alpha)]:\pi_{l}(\alpha(s))\cap[l(x_{1}),l(x_{2})]\neq\emptyset\Big\}.$
In order to show that $\tilde{s}$ is well-defined, it is sufficient to observe that the set appearing in its definition is:
@@
We claim that $\tilde{s}<\mathrm{L}(\alpha)$. If this is not the case, then, combining with (23),
$\pi_{l}\big(\alpha(\mathrm{L}(\alpha))\big)\cap[l(x_{1}+3D),l(x_{2}-3D)]\neq\emptyset.$
Since $X$ is $(R,D)$-thin, $\alpha(\mathrm{L}(\alpha))$ is $D$-close to $l(x_{1}+3D),l(x_{2}-3D)]$, so that
$\pi_{l}\big(\alpha(\mathrm{L}(\alpha))\big)\cap l\big([0,R]\cup[\mathrm{L}(l)-R,\mathrm{L}(l)]\big)=\emptyset,$
a contradiction. This shows that $\tilde{s}<\mathrm{L}(\alpha)$.
Hence, for every $\epsilon>0$, there is $p_{\epsilon}\in\pi_{l}\big(\alpha(\tilde{s}+\epsilon)\big)$. Let $p\in\pi_{l}\big(\alpha(\tilde{s})\big)\cap[l(x_{1}),l(x_{2})]$. Since $X$ is $(R,D)$-thin, it holds $\mathsf{d}(p,\alpha(\tilde{s}))<D$. It follows that $\mathsf{d}(p_{\epsilon},\alpha(\tilde{s}+\epsilon))<D+\epsilon$. Combining these facts, it holds
@@
We now consider item 3. Since the image of $\alpha$ is compact, $\pi_{l}(\alpha)$ is a bounded subset of $l$. Hence, $\pi_{l}(\alpha)\subset[r(0),r(T)]$ for some $T>0$. We consider now the segment $l:=[r(0),r(T+2R)]$. It is easy to check that $\pi_{l}(\alpha)=\pi_{r}(\alpha)$. By item 1, $\pi_{l}(\alpha)$ contains a $3D$-net of either $[l(R),l(t)]$ or $l(t),l(\mathrm{L}(l)-R)]$. This last case cannot happen as $l(t),l(\mathrm{L}(l)-R)]=[r(t),r(T+R)]$ and $\pi_{l}(\alpha)\subset[r(0),r(T)]$. Hence, $\pi_{l}(\alpha)$ (which again coincides with $\pi_{r}(\alpha)$) contains a $3D$-net of $[l(R),l(t)]=[r(R),r(t)]$. ∎
@@
The next proposition shows that an $(R,D)$-thin non-compact metric space is contained in a neighbourhood of controlled width of a ray or a line. We recall that a similar result is proved in [34, Section 5]. The main difference is that Proposition 4.10 below also gives an explicit bound on the width of the aforementioned neighbourhood.
@@
We now prove the statement of the lemma. If $X$ is contained in the $10R$-neighbourhood of $r$, there is nothing to prove. Otherwise, we claim that there exists an element $y\in X$ at maximal distance from $r$. Let $y_{i}\in X$ be a sequence maximizing the distance from $r$. We can assume that $9R\leq\mathsf{d}(y_{i},r)\leq D^{\prime}(X)$. Since $X$ is $(R,D)$-thin, $\pi_{r}(y_{i})\subset[r(0),r(R)]$. It follows that the sequence $\{y_{i}\}_{i\in\mathbb{N}}$ is precompact, so that there exists a limit point $y\in X$ at maximal distance from $r$.
Let $r_{y}$ be the ray which is the limit of the segments joining $y$ and $r(t_{i})$ for a sequence $t_{i}\uparrow+\infty$. By Lemma 4.9, for every $t\geq 0$, it holds $\mathsf{d}(r(t),r_{y})\leq 2R$. Since $\mathsf{d}(y,r)\geq 10R$, it follows that $\pi_{r_{y}}(r)\subset[r_{y}(5R),r_{y}(+\infty))$.
Assume now by contradiction that there exists $y^{\prime}\in X$ such that $\mathsf{d}(y^{\prime},r_{y})\geq 10R$. Let $\gamma$ be a segment realizing the distance between $y^{\prime}$ and $r$. Let $r(t_{\gamma})$ be the endpoint of $\gamma$. Let $r_{y}(t_{\gamma}^{y})\in\pi_{r_{y}}\big(r(t_{\gamma})\big)$. By our previous remarks, $t_{\gamma}^{y}\geq 5R$. By Lemma 4.5, $\pi_{r_{y}}(\gamma)$ contains a $3D$-neighbourhood of $r_{y}(R),r_{y}(5R)]$. Hence, $\gamma$ contains a point $p_{1}$ which is $7D$-close to $r_{y}(R)$. Recall that we also have $\mathsf{d}(y^{\prime},r_{y})\geq 10R$. Hence,
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space. Let $l$ be a segment in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $\bar{l}$ be a segment starting from $l(0)$ with $\pi_{l}(\bar{l}(\mathrm{L}(\bar{l})))\subset[l(0),l(R)]\cup[l(\mathrm{L}(l)-R),l(\mathrm{L}(l)]$. If $\pi_{l}\big(\bar{l}\big)\cap[l(3R),l(\mathrm{L}(l)-3R)]\neq\emptyset$, then $\pi_{l}(\bar{l})$ contains a $3D$-net of $l(R),l(\mathrm{L}(l)-R)]$.
@@
Let $t_{0}\in[0,\mathrm{L}(\bar{l})]$ be such that $\pi_{l}\big(\bar{l}(t_{0})\big)\ni l(t)$ for $t\in[3R,\mathrm{L}(l)-3R]$. Let $\bar{l}_{1}:=\bar{l}_{|[0,t_{0}]}$ and $\bar{l}_{2}:=\bar{l}_{|[t_{0},\mathrm{L}(\bar{l})]}$. Hence, both $\bar{l}_{1}$ and $\bar{l}_{2}$ have an endpoint whose projection on $l$ lies in $l(0),l(R)]\cup[l(\mathrm{L}(l)-R),l(\mathrm{L}(l)]$.
By item 1 of Lemma 4.5, $\pi_{l}(\bar{l}_{1})$ contains a $3D$-net of either $[l(R),l(t)]$ or $l(t),l(\mathrm{L}(l)-R)]$, and the same holds for $\pi_{l}(\bar{l}_{2})$.
@@
Since $\bar{l}(t_{0})$ is $D$-close to a point $l(t)$ in $l(3R),l(\mathrm{L}(l)-3R)]$, it follows
@@
By the same argument, it is not possible that both $\pi_{l}(\bar{l}_{1})$ and $\pi_{l}(\bar{l}_{2})$ contain a $3D$-net of $l(t),l(\mathrm{L}(l)-R)]$, concluding the proof. ∎
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space. Let $l$ be a segment in $X$ and assume that $\mathrm{length}(l)\geq 10R$. Let $p\in X$ be a point with $\mathsf{d}(p,l)\geq 10R$ and let $\bar{l}$ be a segment connecting $p$ to $l(0)$. If $\pi_{l}\big(\bar{l}\big)\cap[l(3R),l(\mathrm{L}(l)-3R)]\neq\emptyset$, then $\mathrm{L}(\bar{l})>\mathrm{L}(l)$.
@@
We now consider the case $b_{1}\in\{\gamma_{1},\gamma_{2}\}$ and $b_{2}=l$. If the statement fails, $\pi_{b_{1}}\big(l\big)\cap(b_{1}(3R),b_{1}(\mathrm{L}(b_{1})-3R))\neq\emptyset$. Consider the orientation of $l$ such that $l(0)\in b_{1}$. Then, $\pi_{b_{1}}(l(\mathrm{L}(l)))\subset[b_{1}(0),b_{1}(R)]\cup[b_{1}(\mathrm{L}(b_{1})-R),b_{1}(\mathrm{L}(b_{1})]$ since otherwise, using that $X$ is $(R,D)$-thin, it follows that $l$ is shorter than $b_{1}$. Hence, by Lemma 4.11, there exists $s>0$ such that $l(s)$ is $R+7D$-close to $p$. This contradicts that $\mathsf{d}(p,l)\geq 10R$.
@@
Assume by contradiction that $\gamma_{1}(t)\in\pi_{\gamma_{1}}(\gamma_{2})$, with $t\in(3R,\mathrm{L}(\gamma_{1})-3R)$. By Lemma 4.11, there exists $s>0$ such that $\gamma_{2}(s)$ is $R+7D$-close to $\gamma_{1}(0)=l(\mathrm{L}(l))$. It follows that
@@
By the previous lemma, $\pi_{b_{2}}\big(b_{1}(0)\big)\cap[b_{2}(3R),b_{2}(\mathrm{L}(b_{2})-3R]\neq\emptyset$. If $\pi_{b_{2}}\big(b_{1}(0)\big)$ intersects $b_{2}(R),b_{2}(3R)]$ or $b_{2}(\mathrm{L}(b_{2})-3R),b_{2}(\mathrm{L}(b_{2})-R)]$, then $\gamma$ has a point which is $5R$-close to either $b_{2}(3R)$ or $b_{2}(\mathrm{L}(b_{2})-3R)$ (because $X$ is $(R,D)$-thin), otherwise the same conclusion follows by Lemma 4.5. ∎